OpenStudy (anonymous):

Your cap size is based on the circumference of your head (in inches). For a certain range of head circumferences c, cap size s can be modeled by the following equation: s=c-1 3 If you wear a size 7 cap, what is your head circumference?

OpenStudy (anonymous):

is that c-1/3?

OpenStudy (anonymous):

c-1 over 3 yes

OpenStudy (anonymous):

ohhh scratch that out then

OpenStudy (anonymous):

oh wow i was not even close i got 15

OpenStudy (anonymous):

okay, this is easy, just substitute 7 to s 7=c-1/3 21=c-1 21+1=c c=22

OpenStudy (anonymous):

ok so your 3(7) to get 21

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay and thats what i had failed to do was substitute 7

OpenStudy (anonymous):

okay pat18 awesome

OpenStudy (anonymous):

pat can u help me pzzz when r done here ")

OpenStudy (anonymous):

yes abalicious?

OpenStudy (anonymous):

pat u still there?